OpenStudy (anonymous):

Solve the equation by graphing: 3x2 + 9x = 12. \ A. {-4, 1} B. {-6, 2} C. {-3, 4} D. No solution

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A i think

OpenStudy (anonymous):

rewrite as 3x^2+9x-12=0 divide through by 3 ( common) +> x^2+3x-4=0 Can solve by either factorization: (x-1)(x+4)=0; hence x =-4 or x =1. You can also use the formula: ax^2+bx+c=0 +> x= [-b (plus or minus)sqrt(b^2-4*a*c)]/{2*a} thus the previous equation yields: x = -24/6 or 6/6 hence x=-4, or x=1

OpenStudy (anonymous):

(x+4)(x-1)= 0 => x = -4 or 1

hero (hero):

3x^2 + 9x - 12 = 0 3x^2 -3x + 12x - 12 = 0 3x(x-1) -12(x-1) = 0 3(x+4)(x-1) = 0

hero (hero):

I acknowledge that I made a mistake. I was in the middle of correcting it however.

hero (hero):

@mas_gh90